# Discrete Spatial Diffusion

### v1

@@ 3.2 Designing Noise Schedules by Structural Similarity Index Metric (SSIM)
The corruption process (1) is a time-homogeneous stochastic process. Consequently, the noise induced in the system, per particle and per unit time, remains constant. However, it has been shown that inhomogeneous noise schedules can facilitate learning (Nichol & Dhariwal, 2021). We use the formulation of a recent study (Santos & Lin, 2023) identified the unique correspondence between non-uniform observation times in a homogeneous Ornstein–Uhlenbeck process (Uhlenbeck & Ornstein, 1930) and noise schedule in conventional diffusion models (Ho et al., 2020; Song et al., 2021a). We follow the same philosophy as Santos & Lin (2023) to construct a sequence of observation times $t_{0}=0<t_{1}<t_{2}<\ldots<t_{T}=1$, at which we will generate random samples for learning. Here, $T$ is the total number of discrete times we will generate corrupted sample images for learning.
We adopt a heuristic approach to construct the discrete times. The idea is to use a metric to quantify how much the “quality” of the images has been degraded up to time $t$, and we aim to design $t_{k}$’s such that the metric degrades from $k=0$ to $k=T$ as evenly as possible. In this manuscript, we chose the Structural Similarity Index Metric (SSIM, Wang et al. (2004)) between the corrupted image and the original one. We generalize a generic monotonic relation between $k$ to $t_{k}$ proposed by Santos et al. (2023):
@@
where $\Phi(p):=\log p/(1-p)$ is the logit function, $\tau_{1}$ and $\tau_{2}$ are parameters used to construct the observation times. Note that $t_{T}=1$ in the above parametrization. Specifically, we tune $\tau_{1}$, $\tau_{2}$ and the unit transition rate $r$ in process (1), using a subset of training samples, aiming to cover an even degradation of the SSIM throughout observation times. We found that setting $\tau_{1}=7.5$ and $\tau_{2}=2.5$, and $r=120$-$160$ is sufficient for numerical experiments. We remark that the choice of the functional form in Eq. (2) is arbitrary and without any theoretical foundation; we only treat Eq. (2) as a versatile monotonic fitting function, whose corresponding SSIM degradation is empirically more symmetric than polynomial and cosine schedules (Nichol & Dhariwal, 2021) for the DSD process (see Appendix Fig. 6).
Figure 2: (a) The forward processes for Gaussian Diffusion (Ho et al., 2020), Inverse Heat Dissipation Model (Rissanen et al., 2022), and Discrete Spatial Diffusion (ours) applied on an image, sampled at discrete times. (b) Percentage change in mass relative to the original image under the forward process.
@@ 3.3 Reverse-time process
Following the general theoretical framework developed in (Campbell et al., 2022; Santos et al., 2023), there exists a reverse-time process that evolves in opposite time and whose joint probability distribution is identical to that of the forward process (1). Specifically, the reverse-time process corresponding to process (1) is
@@
We note that the reverse transition rate depends on both the initial condition $(x_{0},y_{0},c_{0})$ of a particle and the forward solution $p_{t}\left(x,y,c|x_{0},y_{0},c_{0}\right)$, $\forall(x,y,c)$. This is analogous to conventional diffusion models, where either the reverse-time drift (Sohl-Dickstein et al., 2015; Ho et al., 2020) or the score function (Song et al., 2021b) formally depend on the initial sample and the solution of the forward process. However, during the inference, the initial particle configuration is not known, and as such, we train an NN to learn the reverse transition rates using samples $I_{t}$ generated from the forward process (1) at $t>0$. Additionally, the particles are indistinguishable, but the rate prescribed in Eq. 3 is per-particle, raising the question: what is the appropriate per-pixel reverse transition rate that the NN ought to model? This question can be answered by performing the survival analysis of the many-particle system in light of the independence of particle motion; see Appendix A for a derivation. Intuitively, this can also be derived by combining the first-reaction method (Gillespie, 1976) and inhomogeneous Poisson process (e.g., see Corbella et al. (2022)). The analysis shows that the reverse transition rate of the first jump of $n=\left[I_{t}\right]_{x,y,c}$ particles is simply the sum of the instantaneous transition rates:
@@
The above prescribes the rate for the first of all the particles (which is $\left[I_{t}\right]_{x,y,c}$) to jump to one of its neighboring pixels. It also prescribes the rate that the NN will model. This rate is still time-dependent through the dependence on the forward solution $p_{t}$, similar to standard continuous-time diffusion models.
@@ 3.4 Loss functions
Our goal is to provide the corrupted images $I_{t}$ at a sampled time $t>0$ to a neural network (NN) and to train it to predict the reverse-time transition rates (3). We denote the NN modeled rates as $r^{\text{NN}}_{\bar{\nu},x,y,c}(I_{t},t)\in\mathbb{R}_{+}^{4\times H\times W\times C}$. The four in the first dimension here corresponds rates for four nearest-neighbor transitions.
There exist two approaches to formulate the loss functions. The first and more common approach adopts a metric and heuristically matches the NN prediction and the ground truth. DDPM (Ho et al., 2020), score-matching (Song et al., 2021b), and flow-matching (Lipman et al., 2022). When predicting rates, we extend these schemes to “rate-matching”, where we minimize the chosen norm of the difference between the predicted and true rates: $\bar{r}^{\textrm{NN}}$ and $\bar{r}$. For example, for using L1, a loss $\mathcal{L}$:
$\mathcal{L}_{\text{L1}}=\mathbb{E}_{I_{t_{k}},k}\left[\text{mean}(\left|\bar{r}^{\text{NN}}-\bar{r}\log\bar{r}^{\text{NN}}\right|)\right].$ (5)
Here, $k\in\left\{1,\ldots T\right\}$ is uniformly sampled, $I_{t_{k}}$ is drawn from the random process (1) at the sampled times, $\bar{r}=\bar{r}_{\bar{\nu},x,y,c}(I_{t},t|I_{0})$ is the theoretically computed reverse-time transition rate (4), $\bar{r}^{\text{NN}}=\bar{r}^{\text{NN}}_{\bar{\nu},x,y,c}(I_{t},t)$ is the NN-predicted reverse-time transition rate, and the mean is over all the indices $(\bar{\nu},x,y,c)$. The second and more principled approach is through minimization of the negative log-likelihood $L$ of the NN-induced process to predict the analytical reverse-time process (Sohl-Dickstein et al., 2015; Campbell et al., 2022; Santos et al., 2023):
$\mathcal{L}_{L}=-\log L=-\mathbb{E}_{I_{t}}\left[\int_{0}^{\infty}\sum\left(\bar{r}^{\text{NN}}-\bar{r}\log\bar{r}^{\text{NN}}\right){\rm d}t\right].$ (6)
@@
where we again take expectation over randomly sampled $t_{k}$ and $I_{k}$. In this study, we experimented with both loss functions and did not discover any noticeable difference, giving evidence that the DSD forward process (1) is not sensitive to the choice of the loss function. This is favorable over the Gaussian diffusion models as discussed in Ho et al. (2020), which used the heuristic approach to improve over Sohl-Dickstein et al. (2015), which adopted the second approach. We focus on learning the transition rates of the reverse-time dynamics, which is distinct from the ratio-matching approach (Sun et al., 2022; Lou et al., 2024) which focuses on learning the probability distribution $p_{t}(\cdot)$, although a similar formulation (“implicit score entropy”) proposed by (Lou et al., 2024) can be regarded as the process likelihood (7) first proposed in Santos et al. (2023). Algorithm 1 describes the DSD training pseudocode.
@@
3.5 Sampling with Adaptive Time Stepping
@@
Once trained, the neural network will predict reverse-time rates (4), given the configuration of system, $I_{t}$, at time $t\geq 0$. Because the reverse rates are time-dependent, we could generate the exact sample paths of the inhomogeneous Poisson process by integrating the survival function of the first reaction on each pixel in each color channel (see e.g., algorithms reported in Corbella et al. (2022)). However, this approach is not computationally efficient, so we resort to $\tau$-leaping (Gillespie, 2001), an integrator that has been adopted by essentially all continuous-time and discrete-state diffusion models (Campbell et al., 2022; Santos et al., 2023; Winkler et al., 2024; Ren et al., 2024), analogous to the Euler’s method for ordinary or partial differential equations and Euler–Maruyama for Itô SDEs. The central idea of $\tau$-leaping is to approximate the reverse-time transition rates $\bar{r}$ as a fixed constant in a small enough window $(s-\tau,s)$, assuming the time-dependent rates change slowly in the period, a condition often termed as the “leap condition” (Gillespie, 2001; Cao et al., 2005). With this assumption, the original $\tau$-leaping algorithm by Gillespie (2001) generates Poisson random numbers to update the system’s discrete states. However, this approach could sometimes lead to a negative population of particles, which cannot happen in the process, due to violations of the leap condition. Mitigation strategies exist (for example, see Gillespie & Petzold (2003), Cao et al. (2005) and Cao et al. (2006)), however, some of them are limited to small reaction networks and not suitable for the DSD sampling task, which involves a very large number of ($4\times H\times W\times C$) of transition rates to estimate.
Figure 3: (a) Unconditional CelebA generation. (b) Unconditional MNIST generation. (c) Unconditional inpainting on MNIST; 15% difference of conditioning intensity between consecutive rows. (d) Conditioned MNIST generations across different intensities.
@@
As such, we propose a more efficient (but arguably cruder) approach to select the stepper $\tau$ adaptively. Our idea is to combine the binomial $\tau$-leaping (Tian & Burrage, 2004; Chatterjee et al., 2005) and the Courant–Friedrichs–Lewy (CFL) condition (Courant et al., 1928) to conservatively determine the adaptive step size $\tau$. Specifically, since the jump scale is fixed at the pixel length scale, the timescale $\tau$ fully determines the CFL condition. The idea is to choose a $\tau$ such that the CFL number is fixed throughout the inference [^2] Even though CFL condition is more commonly used in PDE integrators, the concept can be applied for our stochastic system. Suppose the reverse-time rate is $\bar{r}$. On average, the particle would move at a timescale $1/\bar{r}$ to one of its neighbors, traveling $\Delta x$. Then, the velocity $c=r\Delta x$. The CFL condition is then $c\Delta t/\Delta x$ where in our scheme $\Delta t$ is the $\tau$; thus, the classical CFL convergence condition translates to the obvious bound of transition probability $\bar{r}\tau<1$. This motivates us to ensure a conservative estimation of $\tau$, but enforcing a small $\bar{r}\tau$ to reduce the error.. To achieve this, we compute the reverse-time transition rates $\bar{r}_{\bar{\nu}}$ for each pixel in each channel, noting that the probability of a particle in that channel will jump to one of its neighboring locations is $\bar{r}_{\bar{\nu}}\tau$. Then, we determine $\tau$ by fixing the largest probability across all the pixels and color channels at a constant. Algorithm 2 describes the DSD inference pseudocode.
@@ 4 Computational Experiments
We employ the Noise Conditional Score Network (NCSN++) (Song et al., 2021b) with two modifications: the final convolutional layer outputs 4 times the number of input channels (e.g., 3 for RGB) to represent four directions (up, down, left, right), and we use a SoftPlus activation function to ensure non-negativity in the predicted rates. The hyperparamters can be found in Appendix E.
@@ 4.1 Image synthesis benchmarks
While the primary motivation for developing DSD is to enable generative modeling under a strict intensity constraint, we first demonstrate the approach on MNIST (LeCun et al., 2010) and CelebA (Liu et al., 2015), demonstrating that the approach can achieve reasonable generative performance for these commonly studied datasets. Unconditionally generated samples are shown in Fig. 3 (a) and (b). Generated samples of the CelebA dataset show that complex patterns including human facial features, lighting, and textures can be captured by DSD.
Next, we explored additional applications of mass conservation for handwritten digits. In Fig. 3 (c), we show results from an in-painting experiment with a fixed mask. In this training, the no-flux boundary conditions were implemented inside the image region, and particles outside of this region were fixed. Throughout the generation process, the disordered particles inside the mask align themselves given the structure outside of the mask. Given the same structure outside of the masked region, we varied the number of particles in the active region, leading to the generation of different digits, as exemplified in the Fig. 3 (c). Additionally, we trained a conditional DSD model that employed the standard class-conditioning (Song et al., 2021b). Figure 3 (d) illustrates the class-conditioned generated images with different total numbers of particles, varying from low, typical, and high total intensities. While these do exhibit some artifacts, DSD surprisingly learns the spatial structure of the digits and generates “Bolder” or “Lighter” digits without saturating the upper bound of the intensity (i.e. 255 for uint8). This would not have been precisely realizable using conventional diffusion models.
@@ 4.2 Subsurface rock microstructures
The microstructure of subsurface rocks governs a wide range of physical processes, including fluid transport, electrical resistivity, and mechanical deformation (Blunt et al., 2013). This originates from connected pores on the nano- and micro-scale, which vary in size, structure, and coordination degree across rock types. High-resolution 3D imaging via X-ray microtomography enables detailed pore-scale reconstructions, but these scans are expensive and limited to sample sizes on the order of millimeters to centimeters (Cnudde & Boone, 2013). While direct imaging of rock microstructure is costly, measuring porosity (defined here as average intensity over the image) across large formations is inexpensive and can be performed without specialized equipment (Leonard, 1948; Passey et al., 1990). This enables large-scale field measurements of porosity, even when high-resolution microstructural data is unavailable.
To overcome this limitation, synthetic models are frequently used to generate representative pore structures for computational physics studies (Øren & Bakke, 2002). However, conventional reconstruction techniques impose strong geometric assumptions that fail to capture the heterogeneity observed in real rocks like Berea Sandstone, Savonnières Carbonate, and Massangis Carbonate. We trained DSD models using these types of rock samples, which represent a broad spectrum of pore structures (including granular, fossiliferous, and dissolution-driven features) across two lithologies: sandstone and carbonate. A description of the training datasets is provided in Appendix F. Figure 4 presents representative outputs from our models trained on 256$\times$256 binary images. The generated samples successfully replicate key statistical properties of the original datasets, including spatial correlation and pore size distribution, both of which are critical for fluid transport. Given that DSD allows for precise control over total porosity, one can generate synthetic microstructures that match the porosity measured in the field, enabling the reconstruction of representative pore-scale samples even in the absence of direct imaging. The model accurately reconstructs microstructural statistics relevant to flow in the subsurface–for details see Appendix F.1.
Figure 5: Generated cathodes with various exactly conditioned phase volume fractions. Carbon binder domain in black, active material particles in gray, electrolyte in white.

### v2

@@ 3.2 Designing Noise Schedules by Structural Similarity Index Metric (SSIM)
Since the corruption process is time-homogeneous (1), the noise applied to each particle remains constant over time. However, it has been shown that inhomogeneous noise schedules can facilitate learning [55]. We use the formulation of a recent study [62] identified the unique correspondence between non-uniform observation times in a homogeneous Ornstein–Uhlenbeck process [73] and noise schedule in conventional diffusion models [36, 67]. We follow the same philosophy as Santos and Lin [62] to construct a sequence of observation times $t_{0}=0<t_{1}<t_{2}<\ldots<t_{T}=1$, at which we will generate random samples for learning. Here, $T$ denotes the total number of discrete time steps used to generate corrupted sample images during training.
We adopt a heuristic approach to construct the discrete times. The idea is to use a metric to quantify how much the “quality” of the images has been degraded up to time $t$, and we aim to design $t_{k}$’s such that the metric degrades from $k=0$ to $k=T$ as evenly as possible. We chose the Structural Similarity Index Metric (SSIM, Wang et al. [76]) between the corrupted image and the original image. We generalize a generic monotonic relation between $k$ to $t_{k}$ proposed by Santos et al. [63]:
@@
where $\Phi(p):=\log p/(1-p)$ is the logit function, $\tau_{1}$ and $\tau_{2}$ are parameters used to construct the observation times. Note that $t_{T}=1$ in the above parametrization. Specifically, we tune $\tau_{1}$, $\tau_{2}$ and the unit transition rate $r$ in process (1), using a subset of training samples, aiming to cover an even degradation of the SSIM throughout observation times. We found that setting $\tau_{1}=7.5$ and $\tau_{2}=2.5$, and $r=120$-$160$ is sufficient for numerical experiments. We remark that the choice of the functional form in Eq. (2) is arbitrary and without any theoretical foundation; we only treat Eq. (2) as a versatile monotonic fitting function, whose corresponding SSIM degradation is empirically more symmetric than polynomial and cosine schedules [55] for the DSD process (see Appendix Fig. 6).
Figure 2: (a) The forward processes for Gaussian Diffusion [36], Inverse Heat Dissipation Model [61], and Discrete Spatial Diffusion (ours) applied on an image, sampled at discrete times. (b) Percentage change in intensity relative to the original image under the forward process.
@@ 3.3 Reverse-time process
Following the process formalism from [11, 63], there exists a reverse-time process that evolves in opposite time and whose joint probability distribution is identical to that of the forward process (1). Specifically, the reverse-time process corresponding to process (1) is:
@@
We note that the reverse transition rate depends on both the initial condition $(x_{0},y_{0},c_{0})$ of a particle and the forward solution $p_{t}\left(x,y,c|x_{0},y_{0},c_{0}\right)$, $\forall(x,y,c)$. This is analogous to conventional diffusion models, where either the reverse-time drift [66, 36] or the score function [69] formally depend on the initial sample and the solution of the forward process. However, during the inference, the initial particle configuration is not known, and as such, we train an NN to learn the reverse transition rates using samples $I_{t}$ generated from the forward process (1) at $t>0$. Additionally, the particles are indistinguishable, but the rate prescribed in Eq. 3 is per-particle, raising the question: what is the appropriate per-pixel reverse transition rate that the NN ought to model? This question can be answered by performing the survival analysis of the many-particle system in light of the independence of particle motion; see Appendix A for a derivation. The analysis shows that the reverse transition rate of the first jump of $n=\left[I_{t}\right]_{x,y,c}$ particles is simply the sum of the instantaneous transition rates:
@@
Eq. 4 prescribes the rate for the first of all the particles (which is $\left[I_{t}\right]_{x,y,c}$) to jump to one of its neighboring pixels. Intuitively, this can also be derived by combining the first-reaction method [30] and inhomogeneous Poisson process (e.g., see Corbella et al. [20]). Eq. 4 also prescribes the rate that the NN will model. This rate is implicitly time-dependent through the dependence on the forward solution $p_{t}$, similar to standard continuous-time diffusion models.
@@ 3.4 Loss functions
Our goal is to provide the corrupted images $I_{t}$ at a sampled time $t>0$ to a neural network (NN) and to train it to predict the reverse-time transition rates $\bar{r}\in\mathbb{R}_{+}^{4\times H\times W\times C}$ of Eq. 4. The first dimension (of size four) represents rates for the four nearest-neighbor transitions. We denote the NN modeled rates as $\bar{r}^{\text{NN}}(I_{t},t)$ with noisified image $I_{t}$ at time $t$.
We use two approaches to formulate the loss function. The first and more commonly used approach adopts a metric to heuristically match the NN prediction and the ground truth, as in DDPM [36], score-matching [69], and flow-matching [49]. We extend these schemes to rate-matching, where we minimize a chosen norm of the difference between the predicted and true rates $\bar{r}^{\textrm{NN}}$ and $\bar{r}$. For example, for using L1, a loss $\mathcal{L}$:
$\mathcal{L}_{\text{rate-matching}}=\mathbb{E}_{I_{t_{k}},k}\left[\text{mean}(\left|\bar{r}^{\text{NN}}-\bar{r}\right|)\right].$ (5)
Here, $k\in\left\{1,\ldots T\right\}$ is uniformly sampled, $I_{t_{k}}$ is drawn from the random process (1) at the sampled times, $\bar{r}=\bar{r}_{\bar{\nu},x,y,c}(I_{t},t|I_{0})$ is the true reverse-time transition rate (4), $\bar{r}^{\text{NN}}=\bar{r}^{\text{NN}}_{\bar{\nu},x,y,c}(I_{t},t)$ is the NN-predicted reverse-time transition rate, and the mean is over all the indices $(\bar{\nu},x,y,c)$. The second and more principled approach is through minimization of the negative log-likelihood $L$ of the NN-induced process to predict the analytical reverse-time process [66, 11, 63]:
$\mathcal{L}_{\text{likelihood}}=-\log L=-\mathbb{E}_{I_{t}}\left[\int_{0}^{\infty}\sum\left(\bar{r}^{\text{NN}}-\bar{r}\log\bar{r}^{\text{NN}}\right){\rm d}t\right].$ (6)
@@
where we again take expectation over randomly sampled $t_{k}$ and $I_{k}$. In this study, we experimented with both loss functions and observed no noticeable difference, giving evidence that the DSD forward process (1) is not sensitive to the choice of the loss function. This stands in contrast to Gaussian diffusion models, where training loss choice significantly impacts performance Ho et al. [36], which used the heuristic approach to improve over Sohl-Dickstein et al. [66], which adopted the second approach. We focus on learning the transition rates of the reverse-time dynamics, which is distinct from the ratio-matching approach [70, 51] which focuses on learning the probability distribution $p_{t}(\cdot)$, although a similar formulation (“implicit score entropy”) proposed by [51] can be regarded as the process likelihood (7) first proposed in Santos et al. [63]. Algorithm 1 describes the DSD training pseudocode.
@@
3.5 Sampling with an Adaptive Time Stepping using the Courant–Friedrichs–Lewy condition
@@
Once trained, the neural network will predict reverse-time rates (4) from the given the configuration of system $I_{t}$, at time $t\geq 0$. Because the reverse rates are time-dependent, one could generate the exact sample paths of the inhomogeneous Poisson process by integrating the survival function of the first reaction on each pixel in each color channel (see e.g., algorithms reported in Corbella et al. [20]). However, this approach is not computationally efficient, so we resort to $\tau$-leaping [31], an integrator that has been adopted by essentially all continuous-time and discrete-state diffusion models [11, 63, 78, 59], analogous to the Euler’s method for ordinary or partial differential equations and Euler–Maruyama for Itô SDEs. The central idea of $\tau$-leaping is to approximate the reverse-time transition rates $\bar{r}$ as a fixed constant in a small enough window $(t-\tau,t)$, assuming the time-dependent rates change slowly in the period, a condition often termed as the “leap condition” [31, 12]. With this assumption, the original $\tau$-leaping algorithm by Gillespie [31] generates Poisson random numbers to update the system’s discrete states. However, this approach could sometimes lead to a negative population of particles, which cannot happen in the process, due to violations of the leap condition. Mitigation strategies exist [32, 12, 13], however, some of them are limited to small reaction networks and not suitable for the DSD sampling task, which involves a very large number of ($4\times H\times W\times C$) of transition rates to estimate.
@@
As such, we propose a more efficient (but arguably cruder) approach to select the step size $\tau$ adaptively. Our idea is to combine the binomial $\tau$-leaping [71, 14] and the Courant–Friedrichs–Lewy (CFL) condition [21] to conservatively determine the adaptive step size $\tau$. Specifically, since the jump scale is fixed at the pixel length scale, the timescale $\tau$ fully determines the CFL condition. The idea is to choose a $\tau$ such that the CFL number is fixed throughout the inference [^2] Even though CFL condition is more commonly used in PDE integrators, the concept can be applied for our stochastic system. Suppose the reverse-time rate is $\bar{r}$. On average, the particle would move at a timescale $1/\bar{r}$ to one of its neighbors, traveling $\Delta x$. Then, the velocity $c=r\Delta x$. The CFL condition is then $c\Delta t/\Delta x$ where in our scheme $\Delta t$ is the $\tau$; thus, the classical CFL convergence condition translates to the obvious bound of transition probability $\bar{r}\tau<1$. This motivates us to ensure a conservative estimation of $\tau$, but enforcing a small $\bar{r}\tau$ to reduce the error.. To achieve this, we compute the reverse-time transition rates $\bar{r}_{\bar{\nu}}$ for each pixel in each channel, noting that the probability of a particle in that channel will jump to one of its neighboring locations is $\bar{r}_{\bar{\nu}}\tau$. Then, we determine $\tau$ by fixing the largest probability across all the pixels and color channels at a constant. Algorithm 2 describes the DSD inference pseudocode.
@@ 4 Computational Experiments
We employ the Noise Conditional Score Network (NCSN++) [69, 1] with two modifications: the final convolutional layer outputs 4 times the number of input channels (e.g., 3 for RGB) to represent four directions (up, down, left, right), and we use a SoftPlus activation function to ensure non-negativity in the predicted rates. The hyperparameters used can be found in Appendix H.
@@ 4.1 Image synthesis benchmarks
While Discrete Spatial Diffusion (DSD) was developed to enable generative modeling under a strict intensity constraint, we first evaluate its performance on conventional image synthesis benchmarks. Specifically, we trained DSD models on MNIST [45], CIFAR-10 [43], and CelebA [50], suggesting that DSD can approximate key statistical and structural properties of image distributions, despite operating under discrete-state and global conservation constraints.
We begin with MNIST, where intensity naturally corresponds to stroke thickness and digit area. In Fig. 3(a), we show inpainting results with a fixed mask with no-flux boundary conditions. During inference, particles inside the masked region rearrange to complete the digit structure, while the surrounding region remains fixed. Holding the visible structure constant, we varied the number of particles allowed to move, revealing that the total intensity governs which digit is most likely to emerge. This highlights the ability of DSD to incorporate hard constraints in downstream tasks such as inpainting. Additionally, we trained a conditional DSD model that employed a standard class-conditioning [69]. In Figure 3(b) we illustrate the class-conditioned generated images with different total numbers of particles, varying from low, typical, to high total intensities. While these do exhibit some artifacts, DSD notably learns the spatial structure of the digits and generates “Bolder” or “Lighter” digits without saturating the upper bound of the intensity (i.e. 255 for uint8). This would not have been precisely realizable using conventional diffusion models. Comparisons with other conditioning approaches can be found in Appendix E.
We also evaluated DSD on RGB datasets to explore scalability and generality. In Fig. 3(c,d), we show unconditional generations from models trained on CIFAR-10 and CelebA. Despite the discrete state space and intensity-preserving dynamics, the model captures complex semantic structures from lighting, and textures to animals, vehicles, and human facial features. Appendix D includes large grids of generated samples, Fréchet Inception Distance (FID), and spatial FID (sFID) metrics, sampling ablations, and post-processing strategies for improving sample smoothness.
Figure 3: (a) Inpainting realizations on MNIST; 15% difference of conditioning intensity between consecutive rows. (b) Conditioned MNIST generations across different intensities and classes. (c) Unconditional CIFAR-10 generations. (d) Unconditional CelebA generations.
@@ 4.2 Subsurface rock microstructures
The microstructure of subsurface rocks governs a wide range of physical processes, including fluid transport, electrical resistivity, and mechanical deformation [7]. This originates from connected pores on the nano- and micro-scale, which vary in size, structure, and coordination degree across rock types. High-resolution 3D imaging via X-ray microtomography enables detailed pore-scale reconstructions, but these scans are expensive and limited to sample sizes on the order of millimeters to centimeters [19]. While direct imaging of rock microstructure is costly, measuring porosity (defined here as average intensity over the image) across large formations is inexpensive and can be performed without specialized equipment [48, 57]. This enables large-scale field measurements of porosity, even when high-resolution microstructural data is unavailable.
Figure 4: Distributions of porosity and generated samples for three rock classes. Each plot shows the porosity distribution across the dataset (shaded curve) and corresponding DSD-generated microstructures at selected quantiles. (Left) Berea Sandstone, (center) Savonnières Carbonate, and (right) Massangis Limestone. Samples illustrate the model’s ability to generate structurally diverse images conditioned on exact global intensity. Quantitative metrics are presented in Appendix F.2.
To overcome this limitation, synthetic models are frequently used to generate representative pore structures for computational physics studies [56]. However, conventional reconstruction techniques impose strong geometric assumptions that fail to capture the heterogeneity observed in real rocks like Berea Sandstone, Savonnières and Massangis Carbonates. We trained DSD models using these rock samples, which represent a broad spectrum of pore structures (including granular, fossiliferous, and dissolution-driven features) across two lithologies. A description of the training datasets is provided in Appendix F.1. Fig. 4 presents representative outputs from our models trained on 256$\times$256 binary images. The generated samples accurately reproduce the key morphological and statistical features of the original datasets, including two-point spatial correlations and pore size distributions which are primary microstructural descriptors that govern permeability, tortuosity, and diffusion behavior in porous media (Fig. 16). Given that DSD allows for precise control over total porosity, one can generate synthetic microstructures that match the porosity measured in the field, enabling the reconstruction of representative pore-scale samples even in the absence of direct imaging.